### v2

# TSA on AutoPilot: Self-tuning Self-supervised Time Series Anomaly Detection

Abstract
Time series anomaly detection (TSAD) finds many applications such as monitoring environmental sensors, industry KPIs, patient biomarkers, etc. A two-fold challenge for TSAD is a versatile and unsupervised model that can detect various different types of time series anomalies (spikes, discontinuities, trend shifts, etc.) without any labeled data. Self-supervised models in particular tackle unsupervised TSAD by transforming the input via various augmentations to create pseudo anomalies for training. However, their performance is sensitive to the choice of augmentation, which is hard to choose in practice, while there exists no effort in the literature on data augmentation tuning for TSAD without labels. Our work aims to fill this gap. We introduce TSAP for TSA “on autoPilot”, which can (self-)tune augmentation hyperparameters end-to-end. It stands on two key components: a differentiable augmentation architecture and an unsupervised validation loss to effectively assess the alignment between augmentation type and anomaly type. Case studies show TSAP’s ability to select the augmentation type and associated hyperparameters.
1 Introduction
Anomaly detection (AD) is a critical task in various domains such as cybersecurity, healthcare, and finance. AD is especially important in time series to ensure system safety and reliability. Thus, there exists a large body of work on time series AD as presented in various surveys [1, 2]. Recent advances in self-supervised learning (SSL) have significantly impacted AD by surpassing traditional unsupervised (or one-class) learning approaches. SSL’s key advantage lies in its ability to self-generate labeled samples, or pseudo anomalies, enabling a more focused exploration of a plausible subspace rather than an exhaustive, impractical search of the entire space. Central to SSL-based AD are data augmentation functions, which create such pseudo labels. These are then used for the (self-)supervised training of an anomaly detector, such as by predicting whether the input is augmented or not [3], which augmentation function is used [4], or through contrastive learning [5]. The success of these approaches highly depends on how well the augmented data mimics the true anomalies [6].
In this paper we introduce TSAP, a novel approach for SSL-based time series anomaly detection (TSAD) “on autoPilot” equipped with end-to-end augmentation hyperparameter tuning. Tuning both discrete and continuous hyperparameters of augmentation enables TSAP to be an anomaly detector that is most suitable for a given task. We summarize our main contributions as follows:
  1. 1.
    Problem: Our work is the first attempt to tune both discrete and continuous hyperparameters of data augmentation in SSL-based TSAD, without labels at training time.
  2. 2.
    New TSAD Method: We propose TSAP  [^1] Code and datasets available at: https://github.com/B-Deforce/TSA-on-autoPilot., which accommodates various time series anomaly types and enables automatic tuning of related hyperparameters (e.g., magnitude, duration) with a differentiable validation loss quantifying alignment between augmented and unlabeled test data.
  3. 3.
    Effectiveness: By carefully selecting augmentation type and its (continuous) hyperparameters, TSAP outperforms existing unsupervised and self-supervised approaches, including the SOTA NeuTraL-AD [7] which also employs learnable augmentations.
2 TSAP: Time Series Anomalies on AutoPilot
Figure 1: Our TSAP framework for end-to-end self-tuning TSAD. Left: Offline trained, differentiable augmentation model $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$ takes as input the normal data and augmentation hyperparameter(s) $\mathbf{a}$, and outputs pseudo-anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$. Right: Self-tuning engine incorporates the pre-trained $f_{\mathrm{aug}}$ (with parameters $\boldsymbol{\phi}$ frozen), alternating between two phases as outlined in Algo. 1.
There are two notable challenges that need to be addressed for automatic selection of both discrete and continuous hyperparameters for SSL-based TSAD:
  1. C1.
    Differentiable Augmentation: Developing an augmentation function that is differentiable with respect to its hyperparameters, enabling gradient-based optimization.
  2. C2.
    Comparable Validation Loss: Formulating a validation loss that quantifies alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{test}}$ while being comparable across different hyperparameter initializations.
To address C1 and C2, we center our framework around two main ideas. First, TSAP employs a differentiable augmentation module implemented as an Encoder-Decoder neural network, $f_{\mathrm{aug}}$, parameterized by $\boldsymbol{\phi}$. This module approximates the anomaly-generating mechanism, conditioned on $\mathbf{a}\in\mathcal{A}^{P}$, $P\geq 1$, where $\mathcal{A}^{P}$ represents the domain of all possible hyperparameter values (e.g., magnitude, duration, …). Importantly, this module is pre-trained independently, establishing it as an offline component of the framework. Second, at test time online, TSAP iteratively refines the detector $f_{\mathrm{det}}$’s parameters $\boldsymbol{\theta}$ as well as augmentation hyperparameters $\mathbf{a}$, through alternating detection and alignment phases. Alignment is performed on part of the unlabeled $\mathcal{D}_{\mathrm{test}}$, referred to as $\mathcal{D}_{\mathrm{val}}$. Based on this, we propose TSAP, a self-tuning self-supervised TSAD framework illustrated in Fig. 1.
2.1 Differentiable Augmentation Module
Anomaly Injection Scheme
We carefully consider accommodating six types of anomalies that are common in real-world time series; namely, trend, extremum, amplitude, mean shift, frequency shift, and platform. We provide a detailed description with visual examples of the anomalies in Appx. A. Each type of anomaly has three hyperparameters; including its starting position (location), duration (length), and severity (level). Extremum captures a spike and only has two hyperparameters as its duration is always $1$. Based on $\mathcal{D}_{\mathrm{trn}}$, the anomaly generation scheme $g$ creates an augmented dataset $\mathcal{D}_{\mathrm{aug}}=\{g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})\mid\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}},\mathbf{a}\sim\mathcal{A}^{P}\}$, where $\mathbf{a}$ is the vector of augmentation hyperparameters that are uniformly randomly sampled from the hyperparameter domain $\mathcal{A}^{P}$.
Model Design
To build an augmentation model $f_{\mathrm{aug}}$ for time series, we use a convolutional neural network (CNN) to encode the input time series $\mathbf{x}_{\mathrm{trn}}$ into the feature map $\mathbf{z}_{\mathrm{trn}}$. We then encode the augmentation hyperparameters $\mathbf{a}$ into $\mathbf{z}_{\mathrm{a}}$, which has the same shape as $\mathbf{z}_{\mathrm{trn}}$, with a multilayer perceptron (MLP). Since the feature map $\mathbf{z}_{\mathrm{trn}}$ generated by the CNN encoder keeps the positional information of the original time series, adding $\mathbf{z}_{\mathrm{a}}$ to $\mathbf{z}_{\mathrm{trn}}$ ensures that only the part with the desired location and length in $\mathbf{z}_{\mathrm{aug}}$ is manipulated. To ensure that the feature maps $\mathbf{z}_{\mathrm{aug}}$ and $\mathbf{z}_{\mathrm{trn}}$ are in the same embedding space, they share the same decoder to reconstruct back to the time series $\mathbf{x}_{\mathrm{aug}}=g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ and $\mathbf{x}_{\mathrm{trn}}$. As such, the loss function of $f_{\mathrm{aug}}$ is based on the reconstruction of both $\mathcal{D}_{\mathrm{trn}}$ and $\mathcal{D}_{\mathrm{aug}}$:
$\mathcal{L}_{\mathrm{aug}}=\sum_{\begin{subarray}{c}\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}}\\ \mathbf{a}\sim\mathcal{A}^{P}\end{subarray}}\big{(}{(\mathbf{x}_{\mathrm{trn}}-\mathbf{\tilde{x}}_{\mathrm{trn}})}^{2}+{(g(\mathbf{x}_{\mathrm{trn}},\mathbf{a})-\mathbf{\tilde{x}}_{\mathrm{aug}})}^{2}\big{)}$ (1)
where $\mathbf{\tilde{x}}_{\mathrm{aug}}$ denotes the output of $f_{\mathrm{aug}}(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ for the same $\mathbf{a}$ as sampled by $g$ (see Fig. 1 left).
2.2 Self-Tuning Module
Algorithm 1 Self-Tuning Module of TSAP
1:Input: Train & val. data $\mathcal{D}_{\mathrm{trn}}$& $\mathcal{D}_{\mathrm{val}}$, pre-trained aug. model $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$, detector $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta})$, max epoch $T$, #inner-loops $L$, and step sizes $\alpha$ and $\beta$
2:Output: Optimized aug. hyperparameters $\mathbf{a}^{*}$ and optimized parameters $\boldsymbol{\theta}^{*}$ corresponding to the lowest $\mathcal{L}_{\mathrm{val}}^{*}$
3:$\mathbf{a}^{(0)}\sim\mathrm{Uniform}(\mathcal{A}^{P})$
4:for $t=0,1,\ldots,T-1$ do
5:   $\boldsymbol{\theta}^{\prime}\leftarrow\boldsymbol{\theta}-\alpha\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\mathrm{trn}}(\boldsymbol{\theta},\mathbf{a}^{(t)})$ for $l$ in $L$triangleright Phase ($i$)
6:   $\mathcal{Z}_{\mathrm{trn}}\leftarrow\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$triangleright Phase ($ii$) starts
7:   $\mathcal{Z}_{\mathrm{aug}}\leftarrow\{f_{\mathrm{det}}^{\mathrm{enc}}(f_{\mathrm{aug}}(\mathbf{x},\mathbf{a}^{(t)}))\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$
8:   $\mathcal{Z}_{\mathrm{val}}\leftarrow\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{val}}\}$
9:   $\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\leftarrow\mathrm{norm.}(\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}})$
10:   $\mathbf{a}^{(t+1)}\leftarrow\mathbf{a}^{(t)}-\beta\nabla_{\mathbf{a}}\mathcal{L}_{\mathrm{val}}(\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}})$
11:   $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}^{\prime}$triangleright Phase ($ii$) ends
12:   if $\mathcal{L}_{\mathrm{val}}^{(t)}<\mathcal{L}_{\mathrm{val}}^{*}$ then $\mathbf{a}^{*}\leftarrow\mathbf{a}^{(t+1)}$;   $\boldsymbol{\theta}^{*}\leftarrow\boldsymbol{\theta}$ end if
13:end for
14:Return $\mathbf{a}^{*}$ and $\boldsymbol{\theta}^{*}$
Central to TSAP is the self-tuning module, which operates by iteratively refining the detector’s parameters, $\boldsymbol{\theta}$, and the augmentation hyperparameters, $\mathbf{a}$. The process is structured into two phases: detection and alignment (see Fig. 1 right). The overall algorithm is given in Algo. 1.
Phase ($i$) Detection
This phase is focused on estimating the parameters $\boldsymbol{\theta}^{(t)}$ of the detector $f_{\mathrm{det}}$ (comprising of an encoder $f_{\mathrm{det}}^{\mathrm{enc}}$ and a discriminator $f_{\mathrm{det}}^{\mathrm{mlp}}$) by minimizing the cross-entropy loss $\mathcal{L}_{\mathrm{trn}}$. This aims to classify between the normal samples $\mathbf{x}_{\mathrm{trn}}$ and the augmented pseudo anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$ by their embeddings $\mathcal{Z}_{\mathrm{trn}}$ and $\mathcal{Z}_{\mathrm{aug}}$, where $\mathcal{Z}_{\mathrm{trn}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ and $\mathcal{Z}_{\mathrm{aug}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(f_{\mathrm{aug}}(\mathbf{x};\mathbf{a}^{(t)}))\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ denote the embeddings of the training data and augmented data, given the current fixed $\mathbf{a}^{(t)}$ at iteration $t$. Note that the parameters $\boldsymbol{\phi}$ of $f_{\mathrm{aug}}$ are frozen throughout this phase.
Phase ($ii$) Alignment
Subsequently, the alignment phase adjusts $\mathbf{a}$ to optimize the unsupervised differentiable validation loss $\mathcal{L}_{\mathrm{val}}$, computed based on the embeddings from the now-updated $f_{\mathrm{det}}^{\mathrm{enc}}$. $\mathcal{L}_{\mathrm{val}}$’s objective is to measure the degree of alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ in the embedding space, as expressed by the Wasserstein distance in Eq. (2), Appx. B.1. Note that the embeddings are normalized to ensure scale invariance before being passed to $\mathcal{L}_{\mathrm{val}}$. This avoids the trivial solution of achieving alignment by setting all embeddings to $\mathbf{0}$ [8]. As the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ in $\mathcal{L}_{\mathrm{val}}$ are obtained through the updated $f_{\mathrm{det}}^{\mathrm{enc}}$, the optimization needs to track the change in $\boldsymbol{\theta}$ caused by the update of $\mathbf{a}$. As such, TSAP employs a second-order optimization process, similar to [8].
Augmentation Type Selection
While Algo. 1 describes continuous hyperparameter tuning, the discrete hyperparameter (i.e. augmentation/anomaly type) selection is done through grid search as the number of anomaly types is finite. Hence, we initialize TSAP for different augmentation types and compare $\mathcal{L}_{\mathrm{val}}$ across types to select the one that yields the best alignment. The idea is that the wrong augmentation type will have poor alignment, while one that reflects the true anomalies in $\mathcal{D}_{\mathrm{test}}$ will result in better alignment, granted proper tuning of the continuous hyperparameters through Algo. 1.
3 Experiments
Table 1: Detection peformance of baselines w.r.t. $F_{1}$ and AUROC on test data across six TSAD tasks (cf. Appx. C.1, Table 3). TSAP has the lowest average rank w.r.t. $F_{1}$ and AUROC, with low std. dev.
Methods PhysioNet A PhysioNet B PhysioNet C PhysioNet D MoCap A MoCap B Avg. Rank
$F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC
OC-SVM 0.182 0.468 0.182 0.472 0.373 0.803 0.393 0.806 1.000 1.000 0.546 0.806 7.2 (3.7) 6.8 (3.3)
LOF 0.999 1.000 0.999 1.000 0.354 0.738 0.358 0.725 0.196 0.506 0.221 0.577 6.8 (3.9) 6.5 (4.4)
ARIMA 0.885 0.960 0.829 0.965 0.991 0.999 0.999 0.999 0.870 0.955 0.225 0.537 4.3 (3.2) 4.7 (3.7)
IF 0.255 0.587 0.232 0.576 0.183 0.402 0.182 0.356 0.864 0.965 0.342 0.758 8.8 (1.7) 8.7 (1.9)
MP 0.812 0.743 0.812 0.744 0.280 0.712 0.284 0.734 1.000 1.000 1.000 1.000 5.0 (3.6) 4.8 (3.4)
EncDec-LSTM 0.190 0.508 0.190 0.508 0.415 0.812 0.442 0.819 0.980 0.999 0.909 0.996 6.5 (2.0) 6.7 (1.9)
SR-CNN 0.965 0.990 0.964 0.998 0.983 0.999 0.991 0.999 0.302 0.700 0.214 0.512 5.2 (4.2) 4.8 (4.5)
USAD 0.183 0.425 0.184 0.428 0.430 0.822 0.409 0.828 1.000 1.000 1.000 1.000 5.5 (4.0) 5.7 (4.5)
NeuTraL-AD 0.211 0.732 0.263 0.679 0.561 0.868 0.526 0.862 1.000 1.000 1.000 1.000 4.2 (2.9) 3.8 (2.5)
TimeGPT 0.327 0.714 0.318 0.711 0.218 0.580 0.217 0.525 0.348 0.743 0.385 0.683 8.0 (1.9) 8.3 (1.6)
TSAP (ours) 1.000 1.000 1.000 1.000 0.973 0.999 0.991 0.998 0.889 0.969 1.000 1.000 2.3 (2.0) 2.2 (2.0)
We evaluate TSAP on six distinct TSAD tasks. Four of these are conducted in a controlled environment, using the 2017 ECG PhysioNet Challenge dataset [9]. The remaining two are natural anomalies, derived from the CMU Motion Capture (MoCap) dataset. In the former, the anomaly types are manually injected in $\mathcal{D}_{\mathrm{test}}$ based on the types discussed in Sec. 2.1. For the latter, the anomaly types in $\mathcal{D}_{\mathrm{test}}$ are unknown and it is the goal of TSAP to find the type that yields best alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ (part of $\mathcal{D}_{\mathrm{test}}$), expressed by $\mathcal{L}_{\mathrm{val}}$. See Appx. C for further details.
Quantitative Results
Table 1 shows the detection results for these six TSAD tasks. TSAP ranks the best overall in both $F_{1}$ and AUROC, showing that detector $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$, trained through the alternating mechanism of TSAP, is able to generalize to unseen, unlabeled anomalies in $\mathcal{D}_{\mathrm{test}}$. While competing methods perform well on certain tasks, they lack consistency across all TSAD tasks. LOF, for instance, thrives on PhysioNet A and B due to its abrupt, manually injected anomalies, but fails on MoCap, where anomalies emerge more gradually as activities transition naturally. SR-CNN and ARIMA show similar behavior. See Appx. B.2B.3 for model/baseline configuration details.
Reconstruction-based methods like EncDec-LSTM and USAD struggle on PhysioNet due to high variability among inliers, yet excel on MoCap, which exhibits consistent, near-periodic patterns. NeuTraL-AD, despite its augmentation-based approach, struggles with PhysioNet’s real-world ECG signal variability, suggesting its augmentation functions lack robustness in noisy conditions. We remark that only TSAP provides robust and consistent performance across all TSAD tasks.
Qualitative Results
Figure 2: Tuning continuous augmentation hyperparameter(s) with TSAP. Top: For Platform anomalies at true level (red dashed line), various initializations converge near the true value (left) by minimizing val. loss (center), which leads to high detection performance (right). Bottom: Multiple continuous hyperparameters (level and length) converge near true values (left) through minimizing val. loss (center), resulting in high AUROC (right). Diverged results correspond to high val. loss, aiding in rejecting low-performance models.
Figure 3: Tuning the discrete hyperparameter (anomaly type) with TSAP. Top: For true Trend anomaly (black), val. loss favors Trend (green) and Mean shift (blue), both achieving high AUROC (middle) and strong resemblance (right), while rejecting Platform (orange) with poor performance. Bottom: For Jump anomalies in MoCap A with unknown type (black), val. loss favors Platform (purple), leading to high AUROC (middle) and similarity to the true anomaly (right), while rejecting Frequency (red) with poor performance.
Key to TSAP’s consistent performance is $\mathcal{L}_{\mathrm{val}}$ through which TSAP automatically learns the augmentation hyperparameters $\mathbf{a}$. Once $\mathbf{a}$ is determined, the task reduces to a supervised learning problem. Next, we show that TSAP not only effectively tunes the continuous augmentation hyperparameters $\mathbf{a}$, but also that $\mathcal{L}_{\mathrm{val}}$ guides the accurate selection of the discrete hyperparameter (i.e. anomaly type). Consider PhysioNet A, where we aim to tune the continuous hyperparameter $a$, i.e. level, of the Platform anomalies present in $\mathcal{D}_{\mathrm{test}}$. That is, the level in $\mathcal{D}_{\mathrm{test}}$ is fixed and tuning aims to estimate its value using TSAP while the other hyperparameters (location, length) are randomized. Fig. 3 (top) shows TSAP’s estimation process for different initializations of $a$. We observe that the initialization for $a\in\{-0.4,0.6\}$ leads to the true $a$$=$$0.2$ (left). Simultaneously, $\mathcal{L}_{\mathrm{val}}$ drops substantially once TSAP has arrived at the true $a$ (center). This is also reflected in the performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ which soars upon estimation of the true $a$ (right). Conversely, the initialization for $a\in\{0.8,-0.8\}$ leads to a high $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$. Indeed, the performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ now suffers from poor alignment. For PhysioNet B, we estimate both level and length while location is randomized. Fig. 3 (bottom) demonstrates TSAP’s ability to accurately estimate the level and length. Further, Fig. 3 (top) showcases TSAP’s ability to perform discrete hyperparameter selection. Here, TSAP has been initialized and trained with three anomaly types (Mean shift, Platform, Trend) on PhysioNet C (true anomaly type: Trend). $\mathcal{L}_{\mathrm{val}}$ indicates a misalignment between the Platform and Trend types (left), also reflected in the AUROC of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ (center). Note how the Mean shift anomaly type has a low $\mathcal{L}_{\mathrm{val}}$ at the end of the training epochs, reflected in the high AUROC on $\mathcal{D}_{\mathrm{val}}$. This shows that the true underlying anomaly type is not necessarily the only type that yields high alignment, and in turn a high-performing detector. In MoCap datasets, where anomaly types are a priori unknown, TSAP is initialized with different augmentation types (Frequency and Platform) to perform discrete hyperparameter selection. Fig. 3 (bottom) highlights its effectiveness as $\mathcal{L}_{\mathrm{val}}$ clearly prefers one type over the other. Indeed, the natural anomalies defined by jumping signals in MoCap A have close resemblance to platform anomalies. See Appx. D for additional results and ablation studies.
4 Conclusion
We introduced TSAP for self-supervised time series anomaly detection, which is the first attempt that automatically (self-)tunes the augmentation hyperparameters on time series data in an unsupervised manner through a differentiable augmentation model and an unsupervised validation loss to help align augmented and test data. While being the first self-tuning SSL solution to TSAD, our work opens avenues for an expanded catalog of anomaly types and extensions to multivariate time series data.
Acknowledgments
Funding: This work was supported by the Research Foundation - Flanders (FWO) [grant number G0C6721N] and [grant number V436123N].
Supplemental Materials
Appendix A Types of Time Series Anomalies
The anomaly generation scheme $g$ accommodates six types of time series anomalies as discussed in Sec. 2.1. Based on $g$, the goal for $f_{\mathrm{aug}}$ is then to learn to inject anomalies into inliers, conditional on the discrete hyperparameter, anomaly type, along with their corresponding hyperparameters as described below. A visual overview of the anomaly types is provided in Fig. 4.
  • Platform: Starting at timestamp location, the values of a duration length in the time series are equal to a constant value level.
  • Mean shift: Starting at timestamp location, a constant value level is added to the values of a duration length in the time series.
  • Amplitude: Starting at timestamp location, a constant value level is multiplied with the values of a duration length in the time series.
  • Trend: Starting at timestamp location, a series of values $at$ is added to the duration length, where $a$ is the level and $t$ is the timestamp in that duration.
  • Extremum (a.k.a. Spike): A large (either positive or negative) value level is assigned to a single timestamp location in the time series.
  • Frequency shift: Starting at phase location, the frequency of the duration with length phases is increased by a constant value level.
Figure 4: Examples of six different types of time series anomalies; (black) original real-world time series, (red) pseudo anomalies generated by $g$. The goal for $f_{\mathrm{aug}}$ is to learn to inject anomalies into inliers, conditional on the discrete (anomaly type) as well as continuous hyperparameters such as level (magnitude), location (start point) and length (see bottom right).
Appendix B Model Configurations
B.1 Validation Loss
We measure the degree of alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ in the embedding space using the Wasserstein distance [10] as described in Sec. 2.2. The Wasserstein distance is a distance measure between probability distributions, of order $p$ between any two marginal distributions $\mu$ and $\nu$, given by:
$W_{p}(\mu,\nu)=\biggl{(}\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(x,y)\sim\gamma}[d(x,y)^{p}]\biggr{)}^{1/p},$ (2)
[…]
B.2 TSAP configuration
The Encoderϕ in $f_{\mathrm{aug}}$ and Encoderθ in $f_{\mathrm{det}}$ are constructed using 1D CNN blocks [12] (transposed 1D CNN for Decoderϕ) for efficient temporal feature extraction. We carefully choose the number of epochs $T$ to allow sufficient time for the convergence of $\mathbf{a}$, with empirical evidence suggesting that $T=100$ typically suffices. For the number of inner-loops $L$, we opt for $L=5$, aligned with [8], such that $f_{\mathrm{det}}$ has adequate time to learn effective discriminative embeddings for $\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{trn}}$. Table 2 provides a comprehensive overview of the configuration details for the different components of TSAP.
Table 2: Configuration details for $f_{\mathrm{aug}}$ and $f_{\mathrm{det}}$ of TSAP
$f_{\mathrm{aug}}$ configuration
Encoderϕ
   Conv Layer 1 {In: 1, Out: 64, Kernel: 100, Stride: 4, ReLU, BatchNorm}
   Conv Layer 2 {In: 64, Out: 64, Kernel: 100, Stride: 4, ReLU}
Decoderϕ
   TransConv Layer 1 {In: 64, Out: 64, Kernel: 100, Stride: 4, ReLU, BatchNorm}
   TransConv Layer 2 {In: 64, Out: 1, Kernel: 100, Stride: 4}
MLPϕ
   MLP Layer 1 {In: 3, Out: 16, ReLU}
   MLP Layer 2 {In: 16, Out: dim$(\mathcal{Z}_{\mathrm{trn}})$, ReLU}
General Parameters
   Batch Size 64
   # Epochs 500
   Optimizer Adam (LR: 0.002)
$f_{\mathrm{det}}$ configuration
Encoderθ
   Conv Layer 1 {In: 1, Out: 32, Kernel: 10, Stride: 2, ReLU, BatchNorm}
   Conv Layer 2 {In: 32, Out: 16, Kernel: 10, Dilation: 2, Stride: 2}
   Conv Layer 3 {In: 16, Out: 8, Kernel: 10, Dilation: 4, Stride: 4}
   Avg Pooling + Flatten {Kernel: 10, Stride: 3}
   Linear Layer {In: 400, Out: 10}
MLPθ
   MLP Layer 1 {In: 10, Out: 1}
General Parameters
   Dropout 0.2
   Batch Size 64
   Warm Start # Epochs $f_{\mathrm{det}}$ 3
   # Epochs 100
   Optimizer for $\mathbf{a}$ Adam (LR: 0.001)
   Optimizer for $f_{\mathrm{det}}$ Adam (LR: 0.002)
   Mixing Rate (cf. Phase ii) 0.15
B.3 Baseline configurations
We compare TSAP with a selection of established baselines, including traditional and deep learning-based methods with demonstrated efficacy in TSAD [13]. The traditional methods consist of different modeling approaches; namely, One-Class Support Vector Machines (OC-SVM) [14]; Local Outlier Factor (LOF) [15]; (ARIMA) [16]; Isolation Forest (IF) [17]; and the Matrix Profile (MP) [18]. On the deep learning side, we benchmark against the Encoder-Decoder LSTM (EncDec-LSTM) [19]; the Spectral Residual Convolutional Neural Network (SR-CNN) [20]; the Unsupervised Anomaly Detection (USAD) for TSAD [21]; and a recent time series foundation model (TimeGPT) [22]. Lastly, we include a state-of-the-art competing method which learns augmentations in the embedding space, called Neural Transformation Learning for (TS)AD (NeuTraL-AD) [7]. This diverse set of baselines allows for a comprehensive analysis across different approaches within the TSAD domain. The details for the baseline configurations are provided below. All models were trained on a single NVIDIA Tesla P100 GPU.
    OC-SVM: We use author-recommended hyperparameters [23].
    LOF: We use author-recommended hyperparameters [15].
    IF: We use author-recommended hyperparameters [17].
    ARIMA: We use AutoARIMA to select hyperparameters [16].
    MP: We use author-recommendations to set the window size $m$ [18]
    EncDec-LSTM: Similar to the original authors [19], we downsample our time series to obtain time series of length approx. equal to 200. For the remaining hyperparameters, we use the original authors’ recommendations.
    SR-CNN: We use author-recommended hyperparameters [20].
    USAD: Similar to the original authors [21], we downsample our time series to obtain time series of length approx. equal to 200. For the remaining hyperparameters, we use the original authors’ recommendations.
    NeuTraL-AD: We use author-recommended hyperparameters [7]. We tune augmentation type separately for each dataset using labeled validation data.
    TimeGPT: We tune the confidence interval [22] for each dataset using labeled validation data.
Appendix C Experimental Setup
C.1 Dataset Details
PhysioNet
The 2017 PhysioNet Challenge dataset [9] comprises a diverse array of real-world 1-lead 300 Hz ECG recordings. We use ECG recordings, each 9 seconds in length with $K=2700$ time-steps and standardized to have a zero mean and standard deviation of one. Injected anomalies represent 10% of the data. We include a total of seven controlled TSAD tasks based on the PhysioNet data as shown in Table 3, with PhysioNet A-D discussed in Sec. 3 and the remainder in Appx. D. For example in PhysioNet A and B, given the anomaly type (Platform), the task is to infer or tune, respectively, the hyperparameter(s) level only and both level and length, while anomaly location is random (hence not tuned). For PhysioNet C and D, the respective tuning tasks are the same but for a different anomaly type (Trend). Finally, Table 4 (top) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in PhysioNet. The hyperparameters location and length are normalized by $K$. Noting that the extremum anomaly always occurs on a single timestamp in the time series, thus length is always $1$.
Table 3: Anomaly profile of different TSAD tasks.
Dataset Type Level Location Length
PhysioNet PhysioNet A Platform Fixed Random Random
PhysioNet B Platform Fixed Random Fixed
PhysioNet C Trend Fixed Random Random
PhysioNet D Trend Fixed Random Fixed
PhysioNet E Mean shift Fixed Random Random
PhysioNet F Mean shift Fixed Random Fixed
PhysioNet G Extremum Fixed Random N/A
MoCap MoCap A Jump Fixed Random Random
MoCap B Run Fixed Random Random
MoCap
The CMU Motion Capture (MoCap) dataset [^2] http://mocap.cs.cmu.edu/ includes signal data from various sensors on subjects’ bodies as they perform different activities (walking, jumping, or running). As we focus on a univariate setting, only the sensor signal on the left femur is used. We consider the walking signal as normal data, and the signals of jumping and running as anomalies. To generate normal signals, we stitch the walking signals by identifying the start and end points of each gait phase and add random noise; whereas to generate anomalous ones, we stitch walking or running signals at a random location in the normal signal. We further add random noises to augment the normal samples in the dataset. Each signal is normalized between $-1$ and $1$ and truncated to length $K=1500$. This yields two distinct TSAD tasks as shown in Table 3. For each, constructed anomalies represent 10% of the data. Different from PhysioNet A–G where we only tune the continuous hyperparameter(s) for the given (discrete) anomaly type, for MoCap A and B, we aim to tune both the unknown anomaly type that corresponds to Jump and Run behavior, respectively, as well as the (continuous) hyperparameter level while location and length take random values. Finally, Table 4 (bottom) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in MoCap. The hyperparameters location and length of platform anomaly are normalized by $K$. The hyperparameters location and length of frequency anomaly denote the starting gait phase and the length of gait phases, respectively.
Table 4: Hyperparameter space for each type of anomaly.
PhysioNet
Anomaly Types location length level
Platform $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
Mean Shift $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
Amplitude $[100..2000]$ $[400..600]$ $\{0.5k+1|k\in[0..10]\}$
Trend $[100..2000]$ $[400..600]$ $\{0.002k-0.01|k\in[0..10]\}$
Extremum/Spike $\{100k|k\in[1..26]\}$$\{1\}$ $\{3k-15|k\in[0..10]\}$
MoCap
Anomaly Types location length level
Platform $[200..800]$ $[100..200]$ $\{0.2k-1|k\in[0..10]\}$
Frequency Shift $[1..3]$ $[1..6]$ $[1..3]$
C.2 Evaluation Metrics
Our method calculates anomaly scores on an entire sequence level $\mathbf{x}$, similar to [7]. This is a different set-up compared to novelty detection in time series which typically operates on a point level. Detection on a sequence level can be especially important to spot long-term patterns (e.g. Trend anomalies). As such, we use the $F_{1}$ score and the Area Under the Receiver Operating Characteristic Curve (AUROC) as key performance metrics to quantify detection capability of anomalous sequences. All results are reported on the unseen $\mathcal{D}_{\mathrm{test}}$. We determine the optimal $F_{1}$ score, by enumerating all possible thresholds, given by the anomaly scores for a given segment $\mathbf{x}$. We then compute the corresponding precision and recall for each threshold and select those that yield the highest $F_{1}$. As such, AUROC provides a balanced view whereas $F_{1}$ shows the optimal scenario. Both metrics range from 0 to 1, with higher values indicating superior performance.
Appendix D Additional Results
In this section, we present additional results for continuous (D.1) and discrete (D.2) augmentation hyperparameter tuning, as well as several ablation studies (D.3).
D.1 Continuous Augmentation Hyperparameter Tuning
In addition to the results on continuous hyperparameter tuning for PhysioNet A and B (see Fig. 3), we demonstrate TSAP’s efficacy in tuning the continuous hyperparameters on five additional TSAD tasks. These include PhysioNet C and D, which feature Trend anomalies, as well as PhysioNet E and F, showcasing Mean shift anomalies, and PhysioNet G, containing Extremum anomalies (cf. Table 3).
PhysioNet C & D
Figure 5: Tuning the continuous augmentation hyperparameter(s) with TSAP for Trend anomalies. Top: Given Trend anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val. loss (center), and leading to high detection AUROC performance (right). Bottom: Multiple continuous hyperparameters, here both level and length are accurately tuned to near true values (left), as guided by minimizing the val. loss (center), achieving high AUROC (right).
The tuning process of the continuous hyperparameters for the Trend anomalies in PhysioNet C and D is shown in Fig. 5. We observe several initializations for $\mathbf{a}$ that arrive closely to the true level for PhysioNet C, as well as to the true level and length for PhysioNet D. In turn, those initializations yield a detector $f_{\mathrm{det}}$ with high performance on $\mathcal{D}_{\mathrm{val}}$. Note how, for PhysioNet C, $\mathcal{L}_{\mathrm{val}}$ is low across the board. This is likely due to the fact that a Trend anomaly with a subtle slope, has similar characteristics to inliers (see e.g. Fig. 3 bottom left). Yet, TSAP effectively assigns a higher validation loss to the initializations that lead to misaligned cases. This shows the effectiveness of our method even in cases where anomalies are subtle.
PhysioNet E & F
Similarly to Trend anomalies, Mean shift anomalies are inherently subtle, especially when the level is close to zero. We show in Fig. 6 how TSAP properly tunes the continuous hyperparameters for level and length in PhysioNet E and F for several initializations.
PhysioNet G
Lastly, Fig. 7 showcases TSAP’s ability to tune the level of the spike in the Extremum anomalies while the location is randomized. Note that Extremum anomalies have no length by definition. The ECG recordings in PhysioNet contain many natural spikes. As such, validation loss is low by default. Nonetheless, TSAP successfully tunes two out of four initializations and reflects – though subtly – this difference in the validation loss. This again leads to a well-tuned $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$ that performs strongly on $\mathcal{D}_{\mathrm{val}}$.
Figure 6: Tuning the continuous augmentation hyperparameter(s) with TSAP for Mean shift anomalies. Top: Given Mean shift anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val. loss (center), and leading to high detection AUROC performance (right). Bottom: Multiple continuous hyperparameters, both level and length are accurately tuned to near true values (left), as guided by minimizing the val. loss (center), achieving high AUROC (right). Mean shift anomalies have low val. loss by default due to the subtle nature of Mean shift anomalies.
Figure 7: Tuning the continuous augmentation hyperparameter(s) with TSAP for Extremum anomalies. Given Extremum anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val. loss (center), and leading to high detection AUROC performance (right). Note how Extremum anomalies have low val. loss by default due to the natural presence of spikes in ECG data.
D.2 Discrete Augmentation Hyperparameter Tuning
We showcased TSAP’s ability to tune the discrete hyperparameter, anomaly type, in Fig. 3 for controlled and natural TSAD tasks. Given the direct applicability and significance of discrete hyperparameter tuning in real-world contexts, we present extended results for discrete hyperparameter tuning.
MoCap B
Fig. 8 shows the discrete hyperparameter tuning for the unknown anomaly type in MoCap B. TSAP was initialized twice: first with $f_{\mathrm{aug}}$ pre-trained for injecting Frequency shift anomalies, and second with $f_{\mathrm{aug}}$ pre-trained for injecting Platform anomalies. The validation loss (left) indicates a strong alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and the unlabeled $\mathcal{D}_{\mathrm{val}}$ when TSAP is initialized with Frequency shift anomalies. This is also reflected in $f_{\mathrm{det}}$’s performance on $\mathcal{D}_{\mathrm{val}}$ (center). Visually, we can indeed confirm that Frequency shift anomalies (right – red) appear to be more similar to the running pattern (right – black) as opposed to Platform anomalies (right – purple).
Figure 8: Tuning discrete hyperparameter (anomaly type) with TSAP. For Run anomalies in MoCap B with unknown type (black), val. loss favors Frequency shift (red) that leads to high AUROC (center) and mimics well true anomaly (right), and effectively rejects inferior type Platform (purple).
D.3 Ablation Studies
We present four ablation studies on the controlled PhysioNet data to support various design strategies of TSAP: validation loss, self-tuning, second-order optimization, and embedding normalization.
Validation Loss
In Fig. 9 (top), we illustrate the level estimation for PhysioNet C under the condition where our Wasserstein-based $\mathcal{L}_{\mathrm{val}}$ is substituted with a point-wise metric, as used in [8]. This comparison shows that a point-wise validation loss tends to favor solutions where the level of the Trend anomaly approximates zero, essentially neutralizing the anomaly. Although this might produce high alignment, it leads to poor $f_{\mathrm{det}}$ performance in $\mathcal{D}_{\mathrm{val}}$ (right). This shows that the distributional characteristics captured by our $\mathcal{L}_{\mathrm{val}}$ are a key contributing factor to the success of TSAP.
Figure 9: Overview of ablation studies. Top: TSAP’s $\mathcal{L}_{\mathrm{val}}$ is replaced by a point-wise val. loss leading to an erroneous estimation of $a$ (left) and poor performance (right). Bottom: TSAP’s self-tuning module is disabled, $a$ is now randomized. Val. loss indicates poor alignment, reflected in poor performance.
Randomization vs. Tuning
In Fig. 9 (bottom), the self-tuning module is disabled for PhysioNet C, where level is instead randomized (along with location, and length). We observe substantially higher $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment. In turn, $f_{\mathrm{det}}$ struggles to detect the unlabeled anomalies in $\mathcal{D}_{\mathrm{val}}$, showing the utility of TSAP’s systematic hyperparameter (self-)tuning over random choice.
Second-order Opt. Ablation
Fig. 10 (top) shows the level-estimation and performance of $f_{\mathrm{det}}$ on PhysioNet C when the second-order optimization is disabled. Note how the estimation process becomes highly unstable when second-order optimization is disabled. In turn, performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ suffers severely.
Embedding Normalization Ablation
We show the level-estimation and performance of $f_{\mathrm{det}}$ on PhysioNet C when the normalization of the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ obtained through $f_{\mathrm{det}}^{\mathrm{enc}}$ is disabled in Fig. 10 (bottom). While $a$ initialized at $0.01$ eventually leads to the correct level, the estimation process is highly volatile compared to when normalization is enabled as shown in Fig. 5 (top).
Figure 10: Overview of additional ablation studies. Top: We disable second-order optimization in TSAP, leading to a highly unstable estimation process of $a$ (left) and poor performance of $f_{\mathrm{det}}$ (right). Bottom: We disable normalization of the embeddings in TSAP. Estimation of $a$ (left) is volatile and does not converge well, in turn, performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ is poor in most cases (right).